# Fast and Robust Visuomotor Riemannian Flow Matching Policy

### v2

@@ Abstract
Diffusion-based visuomotor policies excel at learning complex robotic tasks by effectively combining visual data with high-dimensional, multi-modal action distributions. However, diffusion models often suffer from slow inference due to costly denoising processes or require complex sequential training arising from recent distilling approaches. This paper introduces Riemannian Flow Matching Policy (RFMP), a model that inherits the easy training and fast inference capabilities of flow matching (FM). Moreover, RFMP inherently incorporates geometric constraints commonly found in realistic robotic applications, as the robot state resides on a Riemannian manifold. To enhance the robustness of RFMP, we propose Stable RFMP (SRFMP), which leverages LaSalle’s invariance principle to equip the dynamics of FM with stability to the support of a target Riemannian distribution. Rigorous evaluation on eight simulated and real-world tasks show that RFMP successfully learns and synthesizes complex sensorimotor policies on Euclidean and Riemannian spaces with efficient training and inference phases, outperforming Diffusion Policies and Consistency Policies.
@@ I Introduction
For instance, DP [2], typically based on a Denoising Diffusion Probabilistic Model (DDPM) [12], requires approximately $100$ denoising steps to generate an action. This translates to roughly $1$ second on a standard GPU. Even faster approaches such as Denoising Diffusion Implicit Models (DDIM) [13], still need $10$ denoising steps, i.e., $0.1$ second, per action [2]. Consistency policy [3] aims to accelerate the inference process by training a student model to mimic a DP teacher with larger denoising steps. Despite providing a more computationally-efficient inference, the CP training requires more computational resources and might be unstable due to the sequential training of the two models. Importantly, training these models becomes more computationally demanding when manipulating data with geometric constraints, e.g., robot end-effector orientations, as the computation of the score function of the diffusion process is not as simple as in the Euclidean case [14]. Furthermore, the inference process also incurs increasing computational complexity.
@@
To overcome these limitations, we propose to learn visuomotor robot skills via a Riemannian flow matching policy (RFMP). Compared to DP, RFMP builds on another kind of generative model: Flow Matching (FM) [15, 16]. Intuitively, FM gradually transforms a simple prior distribution into a complex target distribution via a vector field, which is represented by a simple function. The beauty of FM lies in its simplicity, as the resulting flow, defined by an ordinary differential equation (ODE), is much easier to train and much faster to evaluate compared to the stochastic differential equations of diffusion models. However, as many visuomotor policies are represented in the robot’s operational space, action representations must include both end-effector position and orientation. Thus, the policy must consider that orientations lie on either the $\mathcal{S}^{3}$ hypersphere or the $\operatorname{SO}(3)$ Lie group, depending on the specific parametrization. To properly handle such data, we leverage Riemannian flow matching (RFM) [16], an extension of flow matching that accounts for the geometric constraints of data lying on Riemannian manifolds. In our previous work [17], we introduced the idea of leveraging flow matching in robot imitation learning and presented RFMP, which capitalizes on the easy training and fast inference of FM methods to learn and synthesize end-effector pose trajectories. However, our initial evaluation was limited to simple proof-of-concept experiments on the LASA dataset [18].
@@ II Related work
As the literature on robot policy learning is arguably vast, we here focus on approaches that design robot policies based on flow-based generative models.
Normalizing Flows are arguably the first broadly-used generative models in robot policy learning [22]. They were commonly employed as diffeomorphisms for learning stable dynamical systems in Euclidean spaces [23, 24, 25], with extensions to Lie groups [26] and Riemannian manifolds [27], similarly addressed in this paper. The main drawback of normalizing flows is their slow training, as the associated ODE needs to be integrated to calculate the log-likelihood of the model. Moreover, none of the aforementioned works learned sensorimotor policies based on visual observations via imitation learning.
Diffusion Models [28] recently became state-of-art in imitation learning because of their ability to learn multi-modal and high-dimensional action distributions. They have been primarily employed to learn motion planners [29], and complex control policies [2, 30, 4]. Recent extensions use 3D visual representations from sparse point clouds [6], and employ equivariant networks for learning policies that, by design, are invariant to changes in scale, rotation, and translation [7]. However, a major drawback of diffusion models is their slow inference process. In [2], DP requires $10$ to $100$ denoising steps, i.e., $0.1$ to $1$ second on a standard GPU, to generate each action. Consistency models (CM) [31] arise as a potential solution to overcome this drawback [3, 32, 33]. CM distills a student model from a pretrained diffusion model (i.e., a teacher), enabling faster inference by establishing direct connections between points along the probability path. Nevertheless, this sequential training process increases the overall complexity and time of the whole training phase.
@@
In our previous work [17], we proposed to leverage RFM to learn sensorimotor robot policies represented by end-effector pose trajectories on Riemannian manifolds. Building on a similar idea, subsequent works have used FM along with an equivariant transformer to learn $\operatorname{SE}(3)$-equivariant policies [42], for multi-support manipulation tasks with a humanoid robot [43], and for point-cloud-based robot imitation learning [44]. In this paper, we build upon our previous work, Riemannian Flow Matching Policy (RFMP), to enable the learning of complex visuomotor policies on Riemannian manifolds. Unlike the aforementioned approaches, our work focuses on providing a fast and robust RFMP inference process. We achieve this by constructing the FM vector field using LaSalle’s invariance principle, which not only enhances inference robustness with stability guarantees but also preserves the easy training and fast inference capabilities of RFMP.
@@ III-B1 Euclidean Flow Matching
$[\psi_{t}]_{*}p_{0}({\bm{x}})=p_{0}(\psi_{t}^{-1}({\bm{x}}))\det\left(\frac{\partial\psi_{t}^{-1}({\bm{x}})}{\partial{\bm{x}}}\right).$ (2)
The flow $\psi_{t}({\bm{x}})$ is defined via a vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ by solving the ODE,
@@
[…]
@@
where ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{0}|{\bm{x}}_{1})$ denotes the conditional flow. Note that optimizing the CFM loss (5) is equivalent to optimizing the FM loss (4) as they have identical gradients [15]. The problem therefore boils down to design a conditional vector field $u_{t}({\bm{x}}_{t}|{\bm{x}}_{1})$ that generates a probability path $p_{t}$ satisfying the boundary conditions $p_{0}=p$, $q=p_{1}$. Intuitively, $u_{t}({\bm{x}}_{t}|{\bm{x}}_{1})$ should move a randomly-sampled point at $t=0$ to a datapoint at $t\!=\!1$. Lipman et al. [15] proposed the Gaussian CFM, which defines a probability path from a zero-mean normal distribution to a Gaussian distribution centered at $\bm{x}_{1}$ via the conditional vector field,
@@ III-B2 Riemannian Flow Matching
Diffusion Policy (DP) [2] is primarily trained based on DDPM [12], which performs iterative denoising from an initial noise sample ${\bm{x}}^{K}$, where $K$ denotes the total number of denoising steps. The denoising process follows,
${\bm{x}}_{k-1}=\alpha_{t}({\bm{x}}_{k}-\gamma\epsilon_{\bm{\theta}}({\bm{x}}_{k},k)+\mathcal{N}(0,\sigma^{2}\bm{I})),$ (9)
@@
Consistency Policy (CP) [3] aims to speed up the inference process of DP by leveraging a consistency model that distills the DP as a teacher model. While CP adopts a similar denoising mechanism as DP, it enhances the process by incorporating both the current denoising step $k$ and the target denoising step $t$ as inputs to the denoising network, formalized as $\epsilon_{\theta}({\bm{x}}^{k},k,t)$. However, CP involves a two-stage training process. First, a DP teacher policy is trained. Next, a student policy is trained to mimic the denoising process of the teacher policy. This approach enables CP to achieve faster and more efficient inference while retaining the performance of its teacher model, at the cost of a more complex training process. In contrast, FM training is simulation-free, and features a single-phase training with a simple loss function.
@@ IV-A Riemannian Flow Matching Policy
In contrast to the action horizon vector, the observation vector ${\bm{o}}$ is not defined on a receding horizon but is constructed by randomly sampling only few observation vectors. Specifically, RFMP follows the sampling strategy proposed in [37] which uses: (1) A reference observation ${\bm{o}}^{s-1}$ at time step $s-1$; (2) A context observation ${\bm{o}}^{c}$ randomly sampled from an observation window with horizon $T_{o}$, i.e., $c$ is uniformly sampled from $[s-T_{o},s-2]$; and (3) The time gap $s-c$ between the context observation and reference observation. Therefore, the observation vector is defined as ${\bm{o}}=[{\bm{o}}^{s-1},{\bm{o}}^{c},s-c]$. Notice that, when $T_{o}=2$, we disregard the time gap and the observation is ${\bm{o}}=[{\bm{o}}^{s-1},{\bm{o}}^{c}]$. The aforementioned strategy leads to the following RFMP loss function,
@@ 1 Training Process
Input: Predefined number of function evaluation $N$, learned vector field $v_{{\bm{\theta}}}$, observation vector ${\bm{o}}$ and prior distribution $p_{0}$.
@@
12       Integration on the learned Riemannian vector field $\bm{\xi}_{t+\Delta t}=\text{Exp}_{\bm{\xi}_{t}}(v_{{\bm{\theta}}}(\bm{\xi}_{t},{\bm{o}})\Delta t)$
13       Update time $t=t+\Delta t$
@@
Algorithm 1 RFMP Training & Inference Process
After training RFMP, the inference process, which essentially queries the policy $\pi_{\bm{\theta}}({\bm{a}}|{\bm{o}})$, boils down to the following four steps: (1) Draw a sample ${\bm{a}}_{0}$ from the prior distribution $p_{0}$; (2) Construct the observation vector ${\bm{o}}$; (3) Employ an off-the-shelf ODE solver to integrate the learned vector field $v_{t}({\bm{a}}|{\bm{o}};{\bm{\theta}})$ from ${\bm{a}}_{0}$ along the time interval $t=[0,1]$, and get the generated action sequence ${\bm{a}}=[{\bm{a}}^{s},...,{\bm{a}}^{s+T_{p}}]$; (4) Execute the first $T_{a}$ actions of the sequence ${\bm{a}}$ with $T_{a}\leq T_{p}$. This last step allows the robot to quickly react to environment changes, while still providing smooth predicted actions.
@@
Figure 2: RFMP performance when extending inference time beyond $t=1$ on Euclidean PushT and Robomimic Square tasks.
@@ IV-B Stable Riemannian Flow Matching Policy
Both CFM and RCFM train and integrate the learned vector field $v_{t}({\bm{x}}|{\bm{\theta}})$ within the interval $[0,1]$. However, they do not guarantee that the flow converges stably to the target distribution at $t=1$. Besides, the associated vector field may even display strongly diverging behaviors when going beyond this upper boundary [20]. The aforementioned issues may arise due to numerical inaccuracies when training or integrating the vector field. To solve this problem, Sprague et al. [20] proposed Stable Autonomous Flow Matching (SFM) [20], which equips the dynamics of FM with stability to the support of the target distribution. Here, we propose to improve RFMP with SFM, which we generalize to the Riemannian case, in order to guarantee that the flow stabilizes to the target policy at $t=1$. Our experiments show that this approach not only enhances RFMP’s robustness but also further reduces inference time.
@@ IV-B1 Stable Euclidean FMP
$\lim\limits_{t\to\inf}\psi({\bm{x}},t)\in\{{\bm{x}}\in\mathcal{X}|\mathcal{L}_{u}H({\bm{x}})=0\},$
@@ IV-B2 Stable Riemannian FMP
1 Training Process
@@
Input: Predefined number of function evaluation $N$, learned vector field $v_{{\bm{\theta}}}$, observation vector ${\bm{o}}$ and prior distribution $p_{0}$.
@@ 17             $\Delta t=\varepsilon\leq\frac{1}{\lambda_{{\bm{x}}}}$
18      Integration on the learned Riemannian vector field $\bm{\xi}_{t+\Delta t}=\text{Exp}_{\bm{\xi}_{t}}(v_{{\bm{\theta}}}(\bm{\xi}_{t},{\bm{o}})\Delta t)$,
19       Update time $t=t+\Delta t$
20       Update iteration $k=k+1$
Algorithm 2 SRFMP Training & Inference Process
@@ 20       Update iteration $k=k+1$
The parameters $\lambda_{\bm{x}}$ and $\lambda_{\tau}$ have the same influence as in the Euclidean case. Notice that the spatial part of the stable flow (21) closely resembles the geodesic flow (8) proposed in [16]. Figures 3 and 4 show examples of learned RFM and SRFM flows at times $t=\{0,1,1.5\}$ on the sphere and symmetric positive-definite (SPD) matrices manifold. The RFM flow diverges from the target distribution at times $t>1$, while the SRFM flow is stable and adheres to the target distribution for $t\geq 1$.

### v3

@@ Abstract
Diffusion-based visuomotor policies excel at learning complex robotic tasks by effectively combining visual data with high-dimensional, multi-modal action distributions. However, diffusion models often suffer from slow inference due to costly denoising processes or require complex sequential training arising from recent distilling approaches. This paper introduces Riemannian Flow Matching Policy (RFMP), a model that inherits the easy training and fast inference capabilities of flow matching (FM). Moreover, RFMP inherently incorporates geometric constraints commonly found in realistic robotic applications, as the robot state resides on a Riemannian manifold. To enhance the robustness of RFMP, we propose Stable RFMP (SRFMP), which leverages LaSalle’s invariance principle to equip the dynamics of FM with stability to the support of a target Riemannian distribution. Rigorous evaluation on ten simulated and real-world tasks show that RFMP successfully learns and synthesizes complex sensorimotor policies on Euclidean and Riemannian spaces with efficient training and inference phases, outperforming Diffusion Policies and Consistency Policies.
@@ I Introduction
For instance, DP [2], typically based on a Denoising Diffusion Probabilistic Model (DDPM) [12], requires approximately $100$ denoising steps to generate an action. This translates to roughly $1$ second on a standard GPU. Even faster approaches such as Denoising Diffusion Implicit Models (DDIM) [13] still need $10$ denoising steps, i.e., $0.1$ second, per action [2]. Consistency policy [3] aims to accelerate the inference process by training a student model to mimic a DP teacher with larger denoising steps. Despite providing a more computationally-efficient inference, the CP training requires more computational resources and might be unstable due to the sequential training of the two models. Importantly, training these models becomes more computationally demanding when manipulating data with geometric constraints, e.g., robot end-effector orientations, as the computation of the score function of the diffusion process is not as simple as in the Euclidean case [14]. Furthermore, the inference process also incurs increasing computational complexity.
@@
To overcome these limitations, we propose to learn visuomotor robot skills via a Riemannian flow matching policy (RFMP). Compared to DP, RFMP builds on another kind of generative model: Flow Matching (FM) [15, 16]. Intuitively, FM gradually transforms a simple prior distribution into a complex target distribution via a vector field, which is represented by a simple function. The beauty of FM lies in its simplicity, as the resulting flow, defined by an ordinary differential equation (ODE), is much easier to train and much faster to evaluate compared to the stochastic differential equations of diffusion models. However, as many visuomotor policies are represented in the robot’s operational space, action representations must include both end-effector position and orientation. Thus, the policy must consider that orientations lie on either the $\mathcal{S}^{3}$ hypersphere or the $\operatorname{SO}(3)$ Lie group, depending on the chosen parametrization. To properly handle such data, we leverage Riemannian flow matching (RFM) [16], an extension of flow matching that accounts for the geometric constraints of data lying on Riemannian manifolds. In our previous work [17], we introduced the idea of leveraging flow matching in robot imitation learning and presented RFMP, which capitalizes on the easy training and fast inference of FM methods to learn and synthesize end-effector pose trajectories. However, our initial evaluation was limited to simple proof-of-concept experiments on the LASA dataset [18].
@@
Figure 2: RFMP performance when extending inference time beyond $t=1$ on the Euclidean PushT and Robomimic Square tasks.
@@ II Related work
As the literature on robot policy learning is vast, we here focus on approaches that design robot policies based on flow-based generative models.
Normalizing Flows are arguably the first broadly-used generative models in robot policy learning [22]. They were commonly employed as diffeomorphisms for learning stable dynamical systems in Euclidean spaces [23, 24, 25], with extensions to Lie groups [26] and Riemannian manifolds [27], similarly addressed in this paper. The main drawback of normalizing flows is their slow training, as the associated ODE needs to be integrated to calculate the model log-likelihood. Moreover, none of the foregoing works learned sensorimotor policies based on visual observations via imitation learning.
Diffusion Models [28] recently became state-of-art in imitation learning due to their ability to learn multi-modal and high-dimensional action distributions. They have been primarily employed to learn motion planners [29], and complex control policies [2, 30, 4]. Recent extensions use 3D visual representations from sparse point clouds [6], and employ equivariant networks for learning policies that, by design, are invariant to changes in scale, rotation, and translation [7]. However, a major drawback of diffusion models is their slow inference process. In [2], DP requires $10$ to $100$ denoising steps, i.e., $0.1$ to $1$ second on a standard GPU, to generate each action. Consistency models (CM) [31] arise as a potential solution to overcome this drawback [3, 32, 33]. CM distills a student model from a pretrained diffusion model (i.e., a teacher), enabling faster inference by establishing direct connections between points along the probability path. Nevertheless, this sequential training process increases the overall complexity and time of the whole training phase.
@@
In our previous work [17], we proposed to leverage RFM to learn sensorimotor robot policies represented by end-effector pose trajectories on Riemannian manifolds. Building on a similar idea, subsequent works have used FM along with an equivariant transformer to learn $\operatorname{SE}(3)$-equivariant policies [42], for multi-support manipulation tasks with a humanoid robot [43], and for robot imitation learning with point cloud observations [44]. In this paper, we build upon our previous work, Riemannian Flow Matching Policy (RFMP) [17], to enable the learning of complex visuomotor policies on Riemannian manifolds. Unlike the aforementioned approaches, our work focuses on providing a fast and robust RFMP inference process. We achieve this by constructing the FM vector field using LaSalle’s invariance principle, which not only enhances inference robustness with stability guarantees but also preserves the easy training and fast inference capabilities of RFMP.
@@ III-B1 Euclidean Flow Matching
$\psi_{t}]_{*}p_{0}({\bm{x}})=p_{0}(\psi_{t}^{-1}({\bm{x}}))\det\left(\frac{\partial\psi_{t}^{-1}({\bm{x}})}{\partial{\bm{x}}}\right).$ (2)
The flow $\psi_{t}$ is defined via a vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$} by solving the ODE,
@@
where $\bm{\theta}$} are the learnable parameters, $t\sim\mathcal{U}[0,1]$, and ${\bm{x}}\sim p_{t}({\bm{x}})$. However, the loss (4) is intractable since we actually do not have prior knowledge about $u_{t}$ and $p_{t}$. Instead, Lipman et al. [15] proposed to learn a conditional vector field $u_{t}({\bm{x}}|{\bm{x}}_{1})$ with ${\bm{x}}_{1}$ as a conditioning variable. This conditional vector field generates the conditional probability density path $p_{t}({\bm{x}}|{\bm{x}}_{1})$, which is related to the marginal probability path via $p_{t}({\bm{x}})=\int p_{t}({\bm{x}}|{\bm{x}}_{1})q({\bm{x}}_{1})d{\bm{x}}_{1}$, with $q$ being the unknown data distribution. After reparametrization, this leads to the tractable conditional flow matching (CFM) loss function,
@@
where ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{0}|{\bm{x}}_{1})$ denotes the conditional flow. Note that optimizing the CFM loss (5) is equivalent to optimizing the FM loss (4) as they have identical gradients [15]. The problem therefore boils down to design a conditional vector field $u_{t}({\bm{x}}_{t}|{\bm{x}}_{1})$ that generates a probability path $p_{t}$ satisfying the boundary conditions $p_{0}=p$, $q=p_{1}$. Intuitively, $u_{t}({\bm{x}}_{t}|{\bm{x}}_{1})$ should move a randomly-sampled point at $t=0$ to a datapoint at t=1t\!=\!1. Lipman et al. [15] proposed the Gaussian CFM, which defines a probability path from a zero-mean normal distribution to a Gaussian distribution centered at $\bm{x}_{1}$ via the conditional vector field,
@@ III-B2 Riemannian Flow Matching
Diffusion Policy (DP) [2] is primarily trained based on DDPM [12], which performs iterative denoising from an initial noise sample ${\bm{x}}_{K}$, where $K$ denotes the total number of denoising steps. The denoising process follows,
${\bm{x}}_{k-1}=\alpha_{t}\big{(}{\bm{x}}_{k}-\gamma\epsilon_{\bm{\theta}}({\bm{x}}_{k},k)+\mathcal{N}(0,\sigma^{2}\bm{I})\big{)},$ (9)
@@
Consistency Policy (CP) [3] aims to speed up the inference process of DP by leveraging a consistency model that distills the DP as a teacher model. While CP adopts a similar denoising mechanism as DP, it enhances the process by incorporating both the current denoising step $k$ and the target denoising step $t$ as inputs to the denoising network, formalized as $\epsilon_{\theta}({\bm{x}}_{k},k,t)$. However, CP involves a two-stage training process. First, a DP teacher policy is trained. Next, a student policy is trained to mimic the denoising process of the teacher policy. This approach enables CP to achieve faster and more efficient inference while retaining the performance of its teacher model, at the cost of a more complex training process. In contrast, FM training is simulation-free, and features a single-phase training with a simple loss function.
@@ IV-A Riemannian Flow Matching Policy
In contrast to the action horizon vector, the observation vector $\bm{o}$} is not defined on a receding horizon but is constructed by randomly sampling only few observation vectors. Specifically, RFMP follows the sampling strategy proposed in [37] which uses: (1) A reference observation ${\bm{o}}^{s-1}$ at time step s1s-1; (2) A context observation ${\bm{o}}^{c}$ randomly sampled from an observation window with horizon $T_{o}$, i.e., $c$ is uniformly sampled from $s-T_{o},s-2]$; and (3) The time gap scs-c between the context observation and reference observation. Therefore, the observation vector is defined as ${\bm{o}}=[{\bm{o}}^{s-1},{\bm{o}}^{c},s-c]$. Notice that, when $T_{o}=2$, we disregard the time gap and the observation is ${\bm{o}}=[{\bm{o}}^{s-1},{\bm{o}}^{c}]$. The aforementioned strategy leads to the following RFMP loss function,
@@ 1 Training
Input: Predefined number of function evaluation NN, learned vector field $v_{{\bm{\theta}}}$, observation vector $\bm{o}$}, prior distribution $p_{0}$.
@@
12   Integrate the learned Riemannian vector field $\bm{\xi}_{t+\Delta t}=\text{Exp}_{\bm{\xi}_{t}}(v_{{\bm{\theta}}}(\bm{\xi}_{t},{\bm{o}})\Delta t)$.
13   Update time $t=t+\Delta t$.
@@
Algorithm 1 RFMP Training & Inference
After training RFMP, the inference process, which essentially queries the policy $\pi_{\bm{\theta}}({\bm{a}}|{\bm{o}})$, boils down to the following four steps: (1) Draw a sample ${\bm{a}}_{0}$ from the prior distribution $p_{0}$; (2) Construct the observation vector $\bm{o}$}; (3) Employ an off-the-shelf ODE solver to integrate the learned vector field $v_{t}({\bm{a}}|{\bm{o}};{\bm{\theta}})$ from ${\bm{a}}_{0}$ along the time interval $t=[0,1]$, and get the generated action sequence ${\bm{a}}=[{\bm{a}}^{s},...,{\bm{a}}^{s+T_{p}}]$; and (4) Execute the first $T_{a}$ actions of the sequence $\bm{a}$} with $T_{a}\leq T_{p}$. This last step allows the robot to quickly react to environment changes, while still providing smooth predicted actions.
@@ IV-B Stable Riemannian Flow Matching Policy
Both CFM and RCFM train and integrate the learned vector field $v_{t}({\bm{x}}|{\bm{\theta}})$ within the interval $t=[0,1]$. However, they do not guarantee that the flow converges stably to the target distribution at $t=1$. Besides, the associated vector field may even display strongly diverging behaviors when going beyond this upper boundary [20]. The aforementioned issues may arise due to numerical inaccuracies when training or integrating the vector field. To solve this problem, Sprague et al. [20] proposed Stable Autonomous Flow Matching (SFM), which equips the dynamics of FM with stability to the support of the target distribution. Here, we propose to improve RFMP with SFM, which we generalize to the Riemannian case, in order to guarantee that the flow stabilizes to the target policy at $t=1$. Our experiments show that this approach not only enhances RFMP’s robustness but also further reduces inference time.
@@ IV-B1 Stable Euclidean FMP
$\lim\limits_{t\to\infty}\psi({\bm{x}},t)\in\{{\bm{x}}\in\mathcal{X}|\mathcal{L}_{u}H({\bm{x}})=0\},$
@@ IV-B2 Stable Riemannian FMP
1 Training
Input: Initial parameters $\bm{\theta}$}, prior and target distributions $p_{0}$, $p_{1}$.
@@ 1 Training
Input: Predefined number of function evaluation NN, learned vector field $v_{{\bm{\theta}}}$, observation vector $\bm{o}$}, prior distribution $p_{0}$.
@@ 17    $\Delta t=\varepsilon\leq\frac{1}{\lambda_{{\bm{x}}}}$
18  Integrate the learned Riemannian vector field $\bm{\xi}_{t+\Delta t}=\text{Exp}_{\bm{\xi}_{t}}(v_{{\bm{\theta}}}(\bm{\xi}_{t},{\bm{o}})\Delta t)$.
19   Update time $t=t+\Delta t$.
20   Update iteration $k=k+1$.
Algorithm 2 SRFMP Training & Inference
@@
The parameters $\lambda_{\bm{x}}$ and $\lambda_{\tau}$ have the same influence as in the Euclidean case. Notice that the spatial part of the stable flow (21) closely resembles the geodesic flow (8) proposed in [16]. Figures 3 and 4 show examples of learned RFM and SRFM flows at times $t=\{0,1,1.5\}$ on the sphere and symmetric positive-definite (SPD) matrices manifold. The RFM flow diverges from the target distribution at time $t>1$, while the SRFM flow is stable and adheres to the target distribution for $t\geq 1$.